OpenStudy (anonymous):

6x squared -15x = 0 can anyone show me the steps to solve

OpenStudy (anonymous):

6x^2 -15x = 0 or, 3x(2x-5)=0 either 3x=0, i.e x=0 or 2x-5=0, i.e. x=5/2 therefore, x= 0 or 5/2